what is the interest if the Principal is $720, the Rate is 5 1/4% and the Time is 3 years?
Answer by Theo(13342)

if the interest is compounded monthly (normal bank practice), then $720.00 would be equal to $847.2093772in 3 years.
if the interest is compounded annually, then $720.00 would be equal to $844.1213401 in 3 years.
if the interest was compounded daily, then $720.00 would be equal to $847.4908701 in 3 years.

if the interest was compounded continuously, then $720.00 would be equal to $847.50

continuous compounding is given by the formula:
F = P * e^(i*y)

annual, monthly, daily compounding is given by the formula:

F = P * (1+(i/c)^(y*c)

P is the present amount
F is the future amount
i is the annual interest rate
y is the number of year
c is the compounding period.
e is the scientific constant of 2.718281828

example:

for monthly compounding, the variables become:

P = $725.00
F = F
i = .0525
y = 3
c = 12

the formula of F = P * (1+(i/c))^(y*c) becomes:

F = $724.00 * (1 + (.0525/12))^(*3*12) which becomes:

F = $724.00 * (1.004375)^36 which becomes:

F = $847.2093772